OpenStudy (anonymous):

Find the quotient: 2/9 ÷ 5/4

OpenStudy (anonymous):

(2/9) / (5/4) = 2/9 * 4/5 = (2*4)/(5*9) = 8/45

OpenStudy (kymber):

To divide fractions, multiply by the reciprocal.